OpenStudy (bananas1234):

Simplify: square root x^14

OpenStudy (campbell_st):

write it in index form roots are fractional indices \[(x^{14})^{\frac{1}{2}}\] now apply the index law for power of a power

OpenStudy (bananas1234):

not sure i know how to do that

OpenStudy (bananas1234):

@campbell_st

OpenStudy (campbell_st):

the index law for power of a power requires you to multiply the powers. \[(x^a)^b = x^{a \times b}\]